OpenStudy (anonymous):

the minimum value of f(x)=x^3/3-x^2-3x on [-2,2]?

OpenStudy (anonymous):

Do i use the Extreme Value Theorem?

OpenStudy (anonymous):

if i do the derivative of f(x) ; its x^2 - x -3 . Then what?

OpenStudy (anonymous):

equate it to 0, and evaluate with respect to interval [-2,2]

OpenStudy (anonymous):

the resulting value of x I mean as you equate f'(x)=0

OpenStudy (anonymous):

so derive again right? so it becomes 2x-1. then?

OpenStudy (anonymous):

no we solve for possible minimum points on interval [-2,2]... f'(x)=0\[x^2-x-3=0\]\[(x+1)(x-2)=0\]therefore our minimum points are -1 and 2... are they in interval [-2,2]?

OpenStudy (anonymous):

[-1,2] is inside [-2,2] so yes? so what next?

OpenStudy (anonymous):

if they are... they are the minimum value of f(x) you'r looking for...

OpenStudy (anonymous):

to find the y-ordinate, plug-in x at f(x)...

OpenStudy (anonymous):

the options given to me, were 3 fractions and 0 and -9. :(

OpenStudy (anonymous):

x value I mean...

OpenStudy (anonymous):

wait is the function\[f(x)=\frac{1}{3}x^3-x^2-3x~~?\]

OpenStudy (anonymous):

i substituted them, so the lesser results of either -1 or 2 is my minimum value?

OpenStudy (anonymous):

yes that is the function.

OpenStudy (anonymous):

because the 1st derivative should be \[f'(x)=x^2-2x-3\]

OpenStudy (anonymous):

oh. my bad.

OpenStudy (anonymous):

so x=3 and x=-1 right?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

yes it is...

OpenStudy (anonymous):

so my result for x=3 is -9 and x=-1 is 5/3 so i pick 5/3 because it is smaller?

OpenStudy (anonymous):

yup.. you're right...

OpenStudy (anonymous):

oh wait, in terms of negativity, then -9 is smaller right?

OpenStudy (anonymous):

take note your interval

OpenStudy (anonymous):

x=3 is outside your interval

OpenStudy (anonymous):

oh! Ok, so the answer must be within the interval. of course. Thank you!

OpenStudy (anonymous):

yup... np